OpenStudy (anonymous):

antiderivative of \[\int\limits_{0}^{1} (x^4 + 5)^3 dx\]

OpenStudy (anonymous):

please help me

OpenStudy (helder_edwin):

expand the integrand.

OpenStudy (anonymous):

what do you mean? substitutions?

zepdrix (zepdrix):

Unfortunately with this one, you don't have a nice clean substitution available. You'll need to expand out the binomial. (Multiply out the cube thingy) :D

OpenStudy (anonymous):

one sec thats not the full problem

OpenStudy (anonymous):

\[\int\limits_{0}^{1}x^3(x^4 + 5)\]

zepdrix (zepdrix):

Oh lol

OpenStudy (anonymous):

lol sorry

zepdrix (zepdrix):

Soooo do you see something + it's derivative? :O I see x^4 and an x^3 term.. hmm looks like we can do a nice U sub again.

OpenStudy (anonymous):

u= x^4 + 5 u= dx^3 1/4 u= x^3 dx

OpenStudy (anonymous):

??

zepdrix (zepdrix):

\[u=x^4+5\]\[\frac{ 1 }{ 4 }du=x^3 dx\]

OpenStudy (anonymous):

yea i forgot the du

OpenStudy (anonymous):

next, is it 1/4 antiderivative U^3 dx = 1/4

OpenStudy (anonymous):

U^4/4 | 6/5 = 6^4 - 5^4/ 16

OpenStudy (anonymous):

is it confusing to you?

zepdrix (zepdrix):

no you need to set up the integral correctly :D slow down a sec replace all of the X's and DX with U's and DUUUUUU you still have a dx for some reason in your integral.

zepdrix (zepdrix):

1/4 antiderivative u^3 du <- this shouldn't be a dx anymore. Otherwise yes it looks ok.

OpenStudy (anonymous):

so apart from that everything else is right?

zepdrix (zepdrix):

yes, it looks like you found your new limits for U, so thats good. and plugged them in correctly. yay.

OpenStudy (anonymous):

thanks i have got two questions left

OpenStudy (anonymous):

\[\int\limits_{0}^{1}e^(4T + 3) dT\]

OpenStudy (anonymous):

the (4T + 3) are exponents it is supposed to be e^(4T + 3) dT

OpenStudy (anonymous):

we have to do some subs again right

zepdrix (zepdrix):

We don't have to, but yes we probably should. :) So let's do that just so you can get the rhythm of what is going on with exponentials.

OpenStudy (anonymous):

u=4T + 3 du= 4dT 1/4du= dT

zepdrix (zepdrix):

yah looks good.

OpenStudy (anonymous):

1/4\[\int\limits_{5}^{7}e^4du=1/4 \]

OpenStudy (anonymous):

e^4| 7/3= e^7- e^3/4

zepdrix (zepdrix):

why do your u's look like 4's? i don't understand :o

OpenStudy (anonymous):

huh

zepdrix (zepdrix):

e^4...?

zepdrix (zepdrix):

and also, i think your limits should be from 3 to 7, if im calculating that correctly :o

OpenStudy (anonymous):

when t=0, u=5 when t= 1, u=7 so its supposed to be e^7- e^5/4 right?

zepdrix (zepdrix):

when t=0 u=4*0+3 u=3 right?

zepdrix (zepdrix):

\[\huge \int\limits_0^1 e^{(4T+3)}dT \rightarrow \qquad \frac{1}{4} \int\limits_3^7 e^u du\]

OpenStudy (anonymous):

your right

OpenStudy (anonymous):

next step would be 1/4 e^4| 7/3 yea?

zepdrix (zepdrix):

what is e^4? why isn't it e^u?

OpenStudy (anonymous):

right, i keep making that mistake

OpenStudy (anonymous):

apart from this e^4 mistake, everything else is fine?

zepdrix (zepdrix):

yah looks good.

OpenStudy (anonymous):

thanks last but not least

OpenStudy (anonymous):

the last response you gave me in this thread, was that suppose to be the final answer? http://openstudy.com/users/lukasviktor#/updates/50a83c05e4b0129a3c9022f6

zepdrix (zepdrix):

no we didn't apply the upper and lower limits on it yet

OpenStudy (anonymous):

can we finish it please